OpenStudy (anonymous):

help please:) Starting from 1.5 miles away, a car drives towards a speed checkpoint and then passes it. The car travels at a constant rate of 53 miles per hour. The distance of the car from the checkpoint is given by d=l1.5-53tl . At what times is the car 0.1 miles from the checkpoint? Calculate your answer in seconds. 95.1s and 108.7s10.2s and 101.9s108.7s and 10.2s95.1s and 10.2s

OpenStudy (anonymous):

|1.5 - 53t| = 0.1 you'll need to solve these two equations: 1.5 - 53t = 0.1 or 1.5 - 53t = -0.1 can you take it from here? notice you'll have two answers... one where he's headed towards the checkpoint and the other as he heads away from it.

OpenStudy (anonymous):

ight thanks

OpenStudy (anonymous):

you want answer verifications? or you good from here...?

OpenStudy (anonymous):

if u hav them they that would be helpful haha

OpenStudy (anonymous):

i'll tell u if your answers are right or not...